# Zero-Shot Learning of Causal Models

### v2

@@ 2 Related Works
There is an increasing emphasis on developing amortized training methods in causality research. The primary benefit with these methods is their ability to incorporate information from a variety of datasets, and learn general algorithms that can infer infer causal knowledge from observations in a zero-shot manner. Lorch et al. (2022) and Ke et al. (2022) were the earliest works to develop amortized structure learning methods, where they propose transformer-based architectures to integrate samples from multiple datasets. They sample datasets from synthetic data generators during training, which enables them to use supervised objectives for structure learning. This also connects with the recent literature on in-context learning of function classes using transformers (Müller et al., 2021; Akyürek et al., 2022; Garg et al., 2022; Von Oswald et al., 2023). These methods were improved by Wu et al. (2024) for OOD generalization, and applied for the challenging task of recovering gene regulatory networks (Ke et al., 2023). Further, recent works have developed amortized techniques for ATE estimation (Nilforoshan et al., 2023; Zhang et al., 2023), model selection for causal discovery (Gupta et al., 2023), partial causal discovery (topological order) (Scetbon et al., 2024), etc. In this work, we tackle the novel task of amortized learning to infer the causal mechanisms of SCMs.
Autoregressive Causal Learning.
While a vast majority of the literature on causal discovery concerns structure learning (Chickering, 2002; Peters et al., 2014; Zheng et al., 2018), recent works on causal autoregressive flows (Khemakhem et al., 2021; Javaloy et al., 2023) focus on SOTA generative modeling techniques for learning the causal generative processes induced by SCMs. Khemakhem et al. (2021) proved a novel connection between SCMs and autoregressive flows, as the mapping from noise variables to observable variables in SCMs is a triangular map given the topological order of the causal graph. While their work restricted the functional relationships to additive and affine flows, this was extended by Javaloy et al. (2023) to more flexible triangular monotonic increasing maps. More recently, Scetbon et al. (2024) proposed to directly model SCMs, viewed as fixed-point problems on the ordered nodes, using transformer-based architectures. While these methods enable efficient learning of SCMs and their generative processes, they all require to train a specific generative model per dataset. In contrast, we present a novel extension of FiP (Scetbon et al., 2024) that allows amortized learning of functional relationships across different instances from the functional class of SCMs.
3 Amortized Causal Learning: Setup & Background
Structural Causal Models. An SCM defines the causal generative process of a set of $d$ endogenous (causal) random variables $\bm{V}=\{X_{1},\cdots,X_{d}\}$, where each causal variable $X_{i}$ is defined as a function of a subset of other causal variables ($\bm{V}\setminus\{X_{i}\}$) and an exogenous noise variable $N_{i}$:
Hence, an SCM describes the data-generation process of ${\bm{X}}:=[X_{1},\cdots,X_{d}]\sim{\mathbb{P}}_{{\bm{X}}}$ from the noise variables ${\bm{N}}:=[N_{1},\cdots,N_{d}]\sim{\mathbb{P}}_{{\bm{N}}}$ via the function ${\bm{F}}:=[F_{1},\cdots,F_{d}]$, and a graph ${\bm{\mathcal{G}}}\in\{0,1\}^{d\times d}$ indicating the parents of each variable $X_{i}$, that is $[{\bm{\mathcal{G}}}]_{i,j}:=1$ if $X_{j}\in\emph{PA}(X_{i})$. ${\bm{\mathcal{G}}}$ is assumed to be a directed and acyclic graph (DAG), and SCMs to be Markovian (independent noise variables), denoted as ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})$
In addition, we only consider additive noise models (ANM), which are SCMs of the form $X_{i}=F_{i}(\emph{PA}(X_{i}))+N_{i}$.
DAG-Attention Mechanism.
@@ 3 Amortized Causal Learning: Setup & Background
In FiP (Scetbon et al., 2024) the authors propose to leverage the transformer architecture to learn SCMs from observations. By reparameterizing an SCM according to a topological ordering induced by its graph, the authors show that any SCM can be reformulated as a fixed-point problem of the form ${\bm{X}}=\bm{H}({\bm{X}},{\bm{N}})$ where $\bm{H}$ admits a simple triangular structure:
@@
where $\textnormal{Jac}_{\bm{x}}\bm{H}$, $\textnormal{Jac}_{\bm{n}}\bm{H}$ denote the Jacobian of $\bm{H}$ w.r.t the first and second variables respectively.
Motivated by this fixed-point reformulation, FiP considers a transformer-based architecture to model the functional relationships of SCMs and propose a new attention mechanism to represent DAGs in a differentiable manner. Recall that the standard attention matrix is defined as:
@@
Amortized Learning of SCMs.
The goal of amortized learning is to learn (meta-learning) to predict the solution of similar instances of the same optimization problem (Gordon et al., 2018). More formally, given some random inputs $\textbf{I}\sim P_{\textbf{I}}$ and an objective function $L(\theta,\textbf{I})$, the goal of amortized learning is to learn a parameterized model $\mathcal{M}_{\phi}$ s.t.
$\mathcal{M}_{\phi}(\textbf{I})\simeq\theta^{*}(\textbf{I}):=\text{argmin}_{\theta\in\Theta}L(\theta,\textbf{I})$
To train such a model, amortized learning requires to have access to various pairs $(\textbf{I},\theta^{*}(\textbf{I}))$ and optimize the parameters $\phi$ of $\mathcal{M}$ in a supervised manner (Garg et al., 2022).
In our work, we aim to develop amortized learning techniques to predict functional mechanisms from datasets (and causal graphs). More formally, in our setting, $\textbf{I}:=(D_{\textbf{X}},\mathcal{G})$ and $\theta^{*}(\textbf{I}):=\textbf{F}$ where F is the true SCM generating $D_{\textbf{X}}$. Hence, our goal is to learn a single model capable of zero-shot inference of generative SCMs at test time.
4 Methodology: Conditional FiP
In this section, we introduce the proposed approach, Cond-FiP, composed of two key components: (1) a dataset encoder that generates dataset embeddings based on observations and causal graphs, and (2) a conditional variant of FiP (Scetbon et al., 2024), designed for zero-shot inference of the SCMs of datasets when conditioned on their embeddings produced by the encoder. We first present our dataset encoder, then introduce cond-FiP, and conclude with details on generating new observational and interventional samples.
4.1 Dataset Encoder
The objective of this section is to develop a method capable of producing efficient latent representations of datasets. To achieve this, we propose to train an encoder that predicts the noise samples from their associated observations given the causal structures using amortized learning. The latent representations obtained by our encoder will be used later as dataset embeddings to infer SCMs in a zero-shot manner.
Training Setting.
We consider the amortized setting where at training time, we have access to empirical representations of $K$ SCMs $\big{(}{\mathcal{S}}({\mathbb{P}}^{(k)}_{{\bm{N}}},{\bm{\mathcal{G}}}^{(k)},{\bm{F}}^{(k)})\big{)}_{k=1}^{K}$ that have been sampled independently according to a distribution of SCMs ${\mathcal{S}}({\mathbb{P}}^{(k)}_{{\bm{N}}},{\bm{\mathcal{G}}}^{(k)},{\bm{F}}^{(k)})\sim{\mathbb{P}}_{{\mathcal{S}}}$. These empirical representations, denoted $(D_{\bm{X}}^{(k)},{\bm{\mathcal{G}}}^{(k)})_{k=1}^{K}$ respectively, contain each $n$ observations $D_{\bm{X}}^{(k)}:=[\bm{X}^{(k)}_{1},\dots,\bm{X}^{(k)}_{n}]^{T}\in\mathbb{R}^{n\times d}$, and the causal graph ${\bm{\mathcal{G}}}^{(k)}\in\{0,1\}^{d\times d}$. At training time, we also require to have access to the associated noise samples $D_{\bm{N}}^{(k)}:=[\bm{N}^{(k)}_{1},\dots,\bm{N}^{(k)}_{n}]^{T}\in\mathbb{R}^{n\times d}$, which play the role of the target variable in our supervised task. For the sake of clarity, we will omit the dependence on $k$ in our notation and assume access to the full distribution of SCMs ${\mathbb{P}}_{{\mathcal{S}}}$. Our goal here is to learn a model that given a dataset of observations $D_{\bm{X}}$ and the causal graph associated ${\bm{\mathcal{G}}}$, recovers the true noise $D_{\bm{N}}$ from which the observations have been generated. By solving this prediction task, we expect the trained model to provide efficient dataset embeddings as detailed below.
Encoder Architecture.
Following (Lorch et al., 2021; Scetbon et al., 2024), we propose to encode datasets using a transformer-based architecture that alternatively attends on both the sample and node dimensions of the input. More specifically, after having embedded the dataset $D_{\bm{X}}$ into a higher dimensional space using a linear operation $L(D_{\bm{X}})\in\mathbb{R}^{n\times d\times d_{h}}$ where $d_{h}$ is the hidden dimension, the encoder $E$ alternates the application of transformer blocks, consisting of a self-attention block followed by an MLP block (Vaswani et al., 2017), where the attention mechanism is applied either across the samples $n$ or the nodes $d$. When attending over samples, the encoder uses standard self-attention as defined in (2) without masking ($\bm{M}=\{0\}^{n\times n}$), however, when the model attends over the nodes, we leverage the knowledge of the causal graph to mask the undesirable relationships between the nodes, that is we set $\bm{M}=+\infty\times(1-{\bm{\mathcal{G}}})$, with the convention that $0\times(+\infty)=0$, in the standard attention (2). Finally, the obtained embeddings $E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})\in\mathbb{R}^{n\times d\times d_{h}}$ are passed to a prediction network $H:{\mathbb{R}}^{n\times d\times d_{h}}\rightarrow{\mathbb{R}}^{n\times d}$ defined as 2-hidden layers MLP which brings back the encoded datasets to their original space.
Training Procedure.
To infer the noise samples in a zero-shot manner, we propose to minimize the mean squared error (MSE) of predicting the target noises $D_{\bm{N}}$ from the input ($D_{\bm{X}},{\bm{\mathcal{G}}}$) over the distribution of SCMs ${\mathbb{P}}_{{\mathcal{S}}}$ available during training:
${\mathbb{E}}_{{\mathcal{S}}\sim{\mathbb{P}}_{{\mathcal{S}}}}||D_{\bm{N}}-H\circ E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})||^{2}_{2}$
Further, as we restrict ourselves to the case of ANMs, we can equivalently reformulate our training objective in order to predict the functional relationships rather than the noise samples. Indeed, recall that for an ANM ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})$, we have by definition that $\bm{F}(\bm{X})=\bm{X}-\bm{N}$. Therefore, by denoting the new targets as $\bm{F}(D_{\bm{X}}):=D_{\bm{X}}-D_{\bm{N}}$, we propose instead to train our encoder to predict the evaluations of the functional relationships over the SCM distribution by minimizing:
${\mathbb{E}}_{{\mathcal{S}}\sim{\mathbb{P}}_{{\mathcal{S}}}}||\bm{F}(D_{\bm{X}})-H\circ E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})||^{2}_{2}$
Please check Section A.1 for more details on the encoder training procedure and justification for why recoering noise is equivalent to learning the inverse SCM.
Inference.
Given a new dataset $D_{\bm{X}}$ and its causal graph ${\bm{\mathcal{G}}}$, the proposed encoder is able to both provide an embedding $E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})\in\mathbb{R}^{n\times d\times d_{h}}$, and evaluate the functional mechanisms $\widehat{F}(D_{\bm{X}}):=H\circ E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})$. However, this model alone is insufficient for generating new data, whether observational or interventional. This task will be addressed by our conditional decoder, which is detailed in the following section.
On the Knowledge of Graphs.
Prior works on amortized causal learning (Lorch et al., 2022; Ke et al., 2022) proposed to learn to predict the causal graphs using only observations $D_{\bm{X}}$ on synthetically generated datasets. These methods justify our setup where the true graphs are provided as part of the input context to the model, as we can use them to infer the causal graphs in a zero-shot manner from observations if the true graphs information was not available.
4.2 Cond-FiP: Conditional Fixed-Point Decoder
In this section, we present our proposed approach for zero-shot infernce the functional mechanisms of SCMs via amortized training using synthetically generated datasets. To do so, we propose to extend the formulation of FiP with dataset embeddings $E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})$ obtained by our trained encoder as conditions to infer the correct functional mechanisms of the associated SCMs. See Figure 1 for a sketch of Cond-FiP.
Training Setting.
Analogous to the encoder training setup, we assume that we have access to a distribution of SCMs ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})\sim{\mathbb{P}}_{{\mathcal{S}}}$ at training time, from which we can extract empirical representations of the form $(D_{\bm{X}},{\bm{\mathcal{G}}})$ containing the observations and the associated causal graphs respectively. Here, we aim at learning a single model ${\mathcal{T}}$ that can do zero-shot inference of the functional mechanisms of an SCM given its empirical representation. Formally, we aim at training ${\mathcal{T}}$ such that for given any dataset $D_{\bm{X}}$ and its associated causal graph ${\bm{\mathcal{G}}}$ obtained from an SCM ${\mathcal{S}}({\mathbb{P}}_{{\bm{N}}},{\bm{\mathcal{G}}},{\bm{F}})\sim{\mathbb{P}}_{{\mathcal{S}}}$, the conditional function $\bm{z}\in\mathbb{R}^{d}\to\mathcal{T}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})\in\mathbb{R}^{d}$ induced by the model approximates the true functional relationship $\bm{F}:\bm{z}\in\mathbb{R}^{d}\to\bm{F}(\bm{z})\in\mathbb{R}^{d}$. We achieve this by enabling FiP to be conditioned on dataset embeddings from our dataset encoder, detailed below.
Decoder Architecture.
The design of our decoder is based on the FiP architecture for fixed-point SCM learning, with two major differences: (1) we use the dataset embeddings obtained from our encoder as a high dimensional codebook to embed the nodes, and (2) we leverage adaptive layer norm operators (Peebles & Xie, 2023) in the transformer blocks of FiP to enable conditional attention mechanisms.
Conditional Embedding.
The key change of our decoder compared to the original FiP is in the embedding of the input. FiP proposes to embed a data point $\bm{z}:=[z_{1},\dots,z_{d}]\in\mathbb{R}^{d}$ into a high dimensional space using a learnable codebook $\bm{C}:=[C_{1},\dots,C_{d}]^{T}\in\mathbb{R}^{d\times d_{h}}$ and positional embedding $\bm{P}:=[P_{1},\dots,P_{d}]^{T}\in\mathbb{R}^{d\times d_{h}}$, from which they define:
By doing so, Scetbon et al. (2024) ensures that the embedded samples admit the same causal structure as the original samples. However, this embedding layer is only adapted if the samples considered are all drawn from the same observational distribution, as the representation of the nodes, that is given by the codebook $\bm{C}$, is fixed. In order to generalize their embedding strategy to the case where multiple SCMs are considered, we consider conditional codebooks and positional embeddings adapted for each dataset. Formally, given a dataset $D_{\bm{X}}$ and a causal graph ${\bm{\mathcal{G}}}$, we propose to define the conditional codebook and positional embedding as
$\bm{C}(D_{\bm{X}},{\bm{\mathcal{G}}}):=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{C}}$
$\bm{P}(D_{\bm{X}},{\bm{\mathcal{G}}}):=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{P}}$
where $\mu(D_{\bm{X}},{\bm{\mathcal{G}}}):=\text{MaxPool}(E(L(D_{\bm{X}}),{\bm{\mathcal{G}}}))\in\mathbb{R}^{d\times d_{h}}$ is obtained by max-pooling w.r.t the sample dimension the dataset embedding $E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})\in\mathbb{R}^{n\times d\times d_{h}}$ produced by our trained encoder, and $W_{\bm{C}},W_{\bm{P}}\in\mathbb{R}^{d_{h}\times d_{h}}$ are learnable parameters. Then we propose to embed any point $\bm{z}\in\mathbb{R}^{d}$ conditionally on $(D_{\bm{X}},{\bm{\mathcal{G}}})$ by considering:
Adaptive Transfomer Block.
Once an input $\bm{z}\in\mathbb{R}^{d}$ has been embedded into a higher dimensional space $\bm{z}_{\text{emb}}\in\mathbb{R}^{d\times d_{h}}$, FiP proposes to model SCMs by simulating the reconstruction of the data from noise. Starting from $\bm{n}_{0}\in\mathbb{R}^{d\times d_{h}}$ a learnable parameter, they propose to update the current noise $L\geq 1$ times by computing:
where $h$ refers to the MLP block, and for clarity, we omit both the layer’s dependence on its parameters and the inclusion of layer normalization in the notation. Note that here the authors consider their DAG-Attention (3) mechanism in order to modelize correctly the root nodes of the SCM. To obtain a conditional formulation of their computational scheme, we propose first to replace the starting noise $\bm{n}_{0}$ by a conditional one w.r.t. $(D_{\bm{X}},{\bm{\mathcal{G}}})$ and defined as $\bm{n}_{0}:=\mu(D_{\bm{X}},{\bm{\mathcal{G}}})W_{\bm{n}_{0}}\in\mathbb{R}^{d\times d_{h}}$, where $W_{\bm{n}_{0}}\in\mathbb{R}^{d_{h}\times d_{h}}$ is a learnable parameter. Additionally, we propose to add adaptive layer normalization operators (Peebles & Xie, 2023) to both attention and MLP blocks, where each scale or shift is obtained by applying a 1 hidden-layer MLP to the condition, which in our case is $\mu(D_{\bm{X}},{\bm{\mathcal{G}}})$.
Projection.
To project back the latent representation of $\bm{z}$ obtained from previous stages, that is $\bm{n}_{L}\in\mathbb{R}^{d\times d_{h}}$, we propose to simply use a linear operation to get $\widehat{\bm{z}}=\bm{n}_{L}W_{\text{out}}\in\mathbb{R}^{d}$, where $W_{\text{out}}\in\mathbb{R}^{d_{h}}$ is learnable. In the following, we denote the overall architecture ${\mathcal{T}}$, which given an input $\bm{z}\in\mathbb{R}^{d}$ and some condition $(D_{\bm{X}},{\bm{\mathcal{G}}})$ outputs $\widehat{\bm{z}}$, that is ${\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})=\widehat{\bm{z}}$. Note that, for simplicity, we omit the dependence of $\widehat{\bm{z}}$ on $(D_{\bm{X}},{\bm{\mathcal{G}}})$ in the notation.
Training Procedure.
Recall that our goal is zero-shot inference of the functional mechanisms of SCMs given their empirical representations. Therefore, to train our model ${\mathcal{T}}$, we propose to minimize the reconstruction error of the true functional mechanisms estimated by our model over the distribution of SCMs ${\mathbb{P}}_{{\mathcal{S}}}$. More precisely, for any SCM ${\mathcal{S}}(\mathbb{P}_{\bm{N}},{\bm{\mathcal{G}}},\bm{F})\sim{\mathbb{P}}_{{\mathcal{S}}}$ and its empirical representation $(D_{\bm{X}},{\bm{\mathcal{G}}})$, we aim at minimizing
$\mathbb{E}_{\bm{z}\sim\mathbb{P}_{\bm{X}}}\|{\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})-\bm{F}(\bm{z})\|_{2}^{2}$ (4)
where $\bm{z}\sim\mathbb{P}_{\bm{X}}$ is chosen independent of the random dataset $D_{\bm{X}}$. Therefore, when integrating over the distribution of SCMs, we obtain the following optimization problem:
$\mathbb{E}_{{\mathcal{S}}\sim{\mathbb{P}}_{{\mathcal{S}}}}\mathbb{E}_{\bm{z}\sim\mathbb{P}_{\bm{X}}}\|{\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})-\bm{F}(\bm{z})\|_{2}^{2}\;.$
To compute (4), we propose to sample $n$ independent samples $\bm{X}^{\prime}_{1},\dots,\bm{X}^{\prime}_{n}$ from $\mathbb{P}_{\bm{X}}$, leading to a new dataset $D_{\bm{X}^{\prime}}$ independent of $D_{\bm{X}}$, from which we obtain the following optimization problem:
Therefore our training objective aims at learning ${\mathcal{T}}$ such that for any given empirical representation $(D_{\bm{X}},{\bm{\mathcal{G}}})$ of an unknown SCM ${\mathcal{S}}(\mathbb{P}_{\bm{N}},{\bm{\mathcal{G}}},\bm{F})\sim{\mathbb{P}}_{{\mathcal{S}}}$, the conditional function induced by our model, that is $\bm{z}\to{\mathcal{T}}(\bm{z},D_{\bm{X}},{\bm{\mathcal{G}}})$, is close to the true functional mechanism $\bm{F}$ in the MSE sense. In the following, we explain in more detail how to use cond-FiP for causal generation and inference tasks.
Remark on ANM assumption.
Though our method relies on the ANM assumption for encoder training (Appendix A.1), we do not need this assumption for decoder training! An interesting future work would be to consider more general dataset encoding to get rid of this assumption.
4.3 Inference with Cond-FiP
We provide a summary of inference procedure with Cond-FiP, additional details in Section A.2.
Observational Generation.
Cond-FiP is capable of generating new data samples: given a random vector noise $\bm{n}\sim\mathbb{P}_{\bm{N}}$, we can estimate the observational sample associated according to an unknown SCM ${\mathcal{S}}(\mathbb{P}_{\bm{N}},{\bm{\mathcal{G}}},\bm{F})\sim{\mathbb{P}}_{{\mathcal{S}}}$ as long as we have access to its empirical representation $(D_{\bm{X}},{\bm{\mathcal{G}}})$. Formally, starting from $\bm{n}_{0}=\bm{n}$, we infer the associated observation by computing for $\ell=1,\dots,d$:
After (at most) $d$ iterations, $\bm{n}_{d}$ corresponds to the observational sample associated to the original noise $\bm{n}$ according to our conditional SCM $\mathcal{T}(\cdot,D_{\bm{X}},{\bm{\mathcal{G}}})$. To sample noise from $\mathbb{P}_{\bm{N}}$, we leverage cond-FiP that can estimates noise samples under the ANM assumption by computing $\widehat{D_{\bm{N}}}:=D_{\bm{X}}-{\mathcal{T}}(D_{\bm{X}},E(L(D_{\bm{X}}),{\bm{\mathcal{G}}})$. From these estimated noise samples, we can efficiently estimate the joint distribution of the noise by computing the inverse cdfs of the marginals as proposed in FiP.
Interventional Generation.
Cond-FiP also enables the estimation of interventions given an empirical representation $(D_{\bm{X}},{\bm{\mathcal{G}}})$ of an unkown SCM ${\mathcal{S}}(\mathbb{P}_{\bm{N}},{\bm{\mathcal{G}}},\bm{F})\sim{\mathbb{P}}_{{\mathcal{S}}}$. To achieve this, we start from a noise sample $\bm{n}$, and we generate the associated intervened sample $\widehat{\bm{z}}^{\text{do}}$ by directly modifying the conditional SCM provided by Cond-FiP. More specifically, we modify in place the SCM obtained by Cond-FiP, leading to its interventional version $\mathcal{T}^{\text{do}}(\cdot,D_{\bm{X}},{\bm{\mathcal{G}}})$. Now, generating an intervened sample can be done by applying the loop defined in (5), starting from $\bm{n}$ and using the intervened SCM $\mathcal{T}^{\text{do}}(\cdot,D_{\bm{X}},{\bm{\mathcal{G}}})$ rather than the original one.
5 Experiments
5.1 Experimental Setup

### v3

@@ 5 Related Works
Amortized inference has gained traction in causality research, particularly for structure learning. Early works by Lorch et al. (2022) and Ke et al. (2022) introduced transformer-based models trained on multiple synthetic datasets using supervised objectives for amortized inference of causal structure. Their approach aligns with recent works on in-context learning of function classes via transformers (Müller et al., 2021; Akyürek et al., 2022; Garg et al., 2022; Von Oswald et al., 2023). Subsequent improvements targeted OOD generalization (Wu et al., 2024), bayesian causal discovery (Dhir et al., 2024), and applications to gene regulatory networks (Ke et al., 2023). Beyond structure learning, amortized methods have been developed for ATE estimation (Nilforoshan et al., 2023; Zhang et al., 2023; Sauter et al., 2025), model selection (Gupta et al., 2023), and partial causal discovery tasks such as learning topological order (Scetbon et al., 2024). However, amortized inference of causal mechanisms in SCMs remains unaddressed, which is the central focus of our work.
Autoregressive Causal Learning. Most causal discovery methods focus first on structure learning (Chickering, 2002; Peters et al., 2014; Zheng et al., 2018), followed by per-node maximum likelihood estimation to recover the causal mechanisms (Blöbaum et al., 2022). In constrast, recent works on causal autoregressive flows (Khemakhem et al., 2021; Geffner et al., 2022; Javaloy et al., 2023) focus on SOTA normalizing flow based generative models to infer causal mechanisms. Further, FiP (Scetbon et al., 2024) modeled SCMs as fixed-point problems over causal (topological) ordering of nodes using transformer-based architectures. These approaches efficiently learn SCMs but require training a separate model per dataset. In this work, we extend FiP to enable amortized inference of causal mechanisms across different SCM instances, removing this limitation.
6 Conclusion
In this work, we propose novel methodology for training a single model for amortized inference of SCMs. Cond-FiP not only generalizes to unseens in-distribution instances, but also to a wide range of OOD instances, including larger graphs, complex noise distributions, and real-world data. To the best of our knowledge, this is the first approach to demonstrate the feasibility of learning causal mechanisms in a reusable, foundational manner—paving the way for a paradigmatic shift towards the assimilation of causal knowledge across datasets.
Limitations. Our training is limited to synthetic additive noise SCMs due to the requirement of true noise variables for learning the dataset encoder. However, the conditional FiP decoder (see Section 3.2) does not rely on this assumption and can be applied to general SCMs given pretrained dataset embeddings. A promising direction for future work is to explore more general encoding schemes, such as self-supervised learning, or design an implicit in-context learning approach to remove the need for dataset embeddings via direct attention over the context (Mittal et al., 2024).
While Cond-FiP generalizes to larger graphs, it does not yet benefit from larger context sizes at inference (Section K.1), suggesting the need to scale both the model and training data for richer contexts. Additionally, although Cond-FiP performs well on generating interventional samples, it doesn’t perform well on counterfactual generation (Section K.2). Future work will explore scaling Cond-FiP to larger problem instances and application for more complex tasks (counterfactual generation) in real-world scenarios.
Acknowledgements
We thanks the members of the Machine Intelligence team at Microsoft Research for helpful discussions. We also thank Sarthak Mittal for their suggestion to benchmark Cond-FiP in the scarce data regime. Further, we thank Moksh Jain for their feedback on the draft. Part of the experiments were enabled by the Digital Research Alliance of Canada (https://alliancecan.ca/en) and Mila cluster (https://docs.mila.quebec/index.html). Divyat Mahajan acknowledges support via FRQNT doctoral scholarship (https://doi.org/10.69777/354785) for his graduate studies.
Impact Statement
We propose novel methodology for amortized inference of causal mechanisms in structural causal models, representing an initial step toward the development of causal foundational models. Integrating causal principles into machine learning has been widely suggested to improve robustness and reliability, an important property for high-stakes domains such as healthcare, policy, and scientific discovery. By advancing core methodology in causal inference, our work may indirectly support the creation of machine learning systems that are more transparent and trustworthy. However, our research currently does not target any societal application, and does not pose foreseeable risks or negative consequences.
Appendix
[…]
Appendix A Additional Details on Cond-FiP
A.1 DAG-Attention Mechanism
@@
In FiP [Scetbon et al., 2024] the authors propose to leverage the transformer architecture to learn SCMs from observations. By reparameterizing an SCM according to a topological ordering induced by its graph, the authors show that any SCM can be reformulated as a fixed-point problem of the form ${\bm{X}}=\bm{H}({\bm{X}},{\bm{N}})$ where $\bm{H}$ admits a simple triangular structure:
@@
where $\textnormal{Jac}_{\bm{x}}\bm{H}$, $\textnormal{Jac}_{\bm{n}}\bm{H}$ denote the Jacobian of $\bm{H}$ w.r.t the first and second variables respectively. Motivated by this fixed-point reformulation, FiP considers a transformer-based architecture to model the functional relationships of SCMs and propose a new attention mechanism to represent DAGs in a differentiable manner. Recall that the standard attention matrix is defined as:
@@
A.2 Details on Encoder Training
Additive Noise Model Assumption.
Our method relies on the ANM assumption only for the training the encoder. This is because we require the encoder to predict the noise from data in order to obtain embeddings, and under the ANM assumption, the mapping from data to noise can be easily expressed as $x\to x-F(x)$ where $F$ is the generative functional mechanism of the generative ANM. However, if we were to consider general SCMs, i.e. of the form $X=F(X,N)$, we would need access to the mapping $x\to F^{-1}(x,\cdot)(x)$ (assuming this function is invertible), which for general functions is not tractable. Also, note that the ANM assumption by default ensures invertibility since the jacobian w.r.t noise is a triangular matrix with nonzero diagonal elements. An interesting future work would be to consider a more general dataset encoding (using self-supervised techniques) that do not require the ANM assumption, but we believe this is out of the scope of this work.
We now provide further details on training the encoder and show how recovering the noise is equivalent to learn the inverse causal generative process. Recall that an SCM is an implicit generative model that, given a noise sample N, generates the corresponding observation according to the following fixed-point equation in X
$\textbf{X}=F(\textbf{X},\textbf{N})$
$\textbf{X}=F_{\text{gen}}(\textbf{N})$
$\textbf{N}=F_{\text{gen}}^{-1}(\textbf{X})$
$F_{\text{gen}}^{-1}(z)=z-f(z)$
A.3 Inference with Cond-FiP
Sample Generation.
$\textbf{x}=\mathcal{T}(\textbf{x},D_{\bm{X}},\mathcal{G})+\textbf{n}$
$\bm{z}_{\ell}=\mathcal{T}(\bm{z}_{\ell-1},D_{\bm{X}},\mathcal{G})+\textbf{n}$
After $d$ iterations we obtain the following,
$\bm{z}_{d}=\mathcal{T}(\bm{z}_{d},D_{\bm{X}},\mathcal{G})+\textbf{n}$
Interventional Prediction.
Appendix B Details on Experiment Setup with AVICI Benchmark
B.1 AVICI Benchmark